# Initialization Is Critical to Whether Transformer fits Composite Function by Inference or Memorizing

### v1

@@ Abstract
Transformers have shown impressive capabilities across various tasks, but their performance on compositional problems remains a topic of debate. In this work, we investigate the mechanisms of how transformers behave on unseen compositional tasks using anchor functions. We discover that the parameter initialization scale plays a critical role in determining whether the model learns inferential solutions, which capture the underlying compositional primitives, or symmetric solutions, which simply memorize mappings without understanding the compositional structure. By analyzing the information flow and vector representations within the model, we reveal the distinct mechanisms underlying these solution types. We further find that inferential solutions exhibit low complexity bias, which we hypothesize is a key factor enabling them to learn individual mappings for single anchors. Building upon our understanding of these mechanisms, we can predict the learning behavior of models with different initialization scales when faced with data of varying inferential complexity. Our findings provide valuable insights into the role of initialization scale in shaping the type of solution learned by transformers and their ability to learn and generalize compositional functions.
@@ 1 Introduction
Large-scale transformers demonstrate unprecedented capabilities (Achiam et al., 2023; Teubner et al., 2023; Choi et al., 2021; Brown et al., 2020), even noted as "sparks of AGI" (Bubeck et al., 2023). However, their performance on compositional tasks, which are considered a key property of human cognition (Marcus, 2003; Smolensky et al., 2022), has long been the subject of intense debate, especially in terms of out-of-distribution (OOD) generalization ability. Furthermore, the mechanisms underlying how large language models accomplish compositional tasks remain unclear. This raises critical open questions about how to faithfully interpret transformers’ capabilities on compositional tasks. Can transformers learn the underlying compositional primitives within the data, or do they merely learn input-output mappings? When transformers produce incorrect outputs on compositional tasks, do their responses follow any discernible patterns or logic?
@@
In this work, we use anchor functions (Zhang et al., 2024) to investigate how transformers generalize to unseen compositional tasks. In our setup, each sequence contains a composite anchor pair, a key item, and noise items unrelated to the output (right side of Fig. 1(a)). The single anchors are specific tokens (i.e., 1, 2, 3, 4), each corresponding to a particular arithmetic operation (left side of Fig. 1(a)). A composite function is defined as the sequential application of two single anchor functions, i.e., the corresponding arithmetic operation, to a key item. For example, given a key item $x$ and a composite anchor pair (1, 2), the composite function would output $(x+5)+1$. As shown in the middle part of Fig. 1(a), during training, 14 out of the 16 possible composite anchor pairs are assigned inferential mappings, meaning the target output is consistent with the result of applying the composite function. One pair (3, 4) is assigned a non-inferential mapping, where the target output deviates from the result of the composite function. The remaining pair (4, 3) is held out as an unseen task. After training, there are three possible solutions the model could learn for the unseen pair (4, 3): an inferential solution, a symmetric solution matching the non-inferential seen pair (3, 4), or other non-inferential solutions (Fig. 1(b)). The central question we aim to answer is: which of these three types of solutions will the model learn for the unseen anchor pair (4, 3)?
@@
Figure 1: Experimental setup and possible solutions and mechanisms for the unseen composite anchor pair (4, 3). (a) Data generation: Left: The single anchors (1, 2, 3, 4) correspond to specific arithmetic operations. Middle: During training, 14 out of the 16 possible composite anchor pairs are assigned inferential mappings, one pair (3, 4) is assigned a non-inferential mapping, and the remaining pair (4, 3) is held out as an unseen task (does not appear in the training set). Right: The input sequences comprise a composite anchor pair, a key item preceding the anchor pair, and other unrelated items. The question mark indicates the output for the unseen anchor pair (4, 3), which depends on the learned solution. (b) Possible solutions and mechanisms for the unseen anchor pair (4, 3): After training, the model may learn an inferential solution, a symmetric solution matching the non-inferential seen pair (3, 4), or other non-inferential solutions not shown in the figure. Two potential mechanisms for task generalization are illustrated: learning the symmetric structure (Mechanism 1) or composing the inferred single anchor mappings (Mechanism 2).
@@
Based on the above experimental setup, we find that the model exhibits two distinct phases of solutions depending on the parameter initialization scale with different model depths. With small initialization scales, the model tends to learn inferential solutions on the unseen composite anchor pair (4, 3). In contrast, with larger initialization scales, the model is more likely to learn symmetric solutions on anchor pair (4, 3) that simply match the output of its symmetric anchor pair (3, 4). These findings suggest that the initialization scale plays a crucial role in determining the type of solution learned by the model.
@@
Building upon our understanding of these mechanisms, we predict and experimentally validate that models with small initialization scales require more training steps to fit data with larger inferential complexity, defined as the number of composite anchor pair groups that do not satisfy the inferential mapping, while models with large initialization scales use similar training steps regardless of inferential complexity.
In summary, our work highlights the crucial role of initialization scale in shaping the solutions learned by transformers on compositional tasks. These insights can inform the design of future transformer models, enabling them to better capture compositional structures and generalize to unseen tasks, ultimately leading to improved performance on compositional problems.
@@ 2 Related Work
Challenges of transformers in compositional tasks.
Recent advancements in LLMs have showcased remarkable capabilities across a wide range of tasks, often surpassing human performance (Wei et al., 2022; Fu et al., 2022). However, despite their impressive performance on single-step reasoning tasks (Srivastava et al., 2022), transformers face significant challenges when it comes to effectively combining multiple steps to solve compositionally complex problems (Hupkes et al., 2018; Csordás et al., 2021, 2022; Yun et al., 2022; Lepori et al., 2023; Okawa et al., 2023; Dziri et al., 2024). Ramesh et al. (2023) show that transformers trained to directly compose capabilities struggle to generalize to out-of-distribution (OOD) tasks. Liu et al. (2022) suggest that shallow transformers learn shortcuts during training, leading to poor OOD generalization. Numerous studies have explored various approaches to address these limitations, such as encouraging explicit reasoning step generation within a single generation (Wei et al., 2022), leveraging LLMs to generate reasoning steps iteratively (Creswell et al., 2022; Creswell and Shanahan, 2022). Despite these efforts, achieving complete mastery in compositional tasks remains a significant challenge for vanilla transformers. Zhang et al. (2024) propose a type of benchmark functions called anchor functions for studying language models. Building upon the compositional task settings in anchor functions, our work investigates the solutions and underlying mechanisms of models under different initialization scales.
Parameter initialization of neural networks.
The parameter initialization of the network is important to determine the final fitting result of the network (Jacot et al., 2018; Arora et al., 2019; Zhang et al., 2019; Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2018; Chizat and Bach, 2018; Sirignano and Spiliopoulos, 2020; Williams et al., 2019; E et al., 2020). Luo et al. (Luo et al., 2021), Zhou et al. (Zhou et al., 2022) mainly identify the linear regime and the condensed regime for two-layer and three-layer wide ReLU NNs, respectively. A series of works suggest that the condensed regime improves the generalization ability of models (Zhang et al., 2023; Zhang and Xu, 2024; Zhang et al., 2021, 2022). Recent studies have also investigated the impact of initialization on the training of large language models (LLMs) (Liu et al., 2020; Zhu et al., 2021; Trockman and Kolter, 2023; Zhang et al., 2019; Huang et al., 2020; Wang et al., 2024). These works primarily focus on how the initialization scale affects training stability and other aspects of LLM performance. In our work, we find that different initialization scales can influence a model’s ability to memorize and infer data on the compositional task.
@@ 3.1 Two-anchor composite function
A two-anchor composite function $f(X):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{d}$ is defined as
$f(\bm{x}_{1},\ldots,\bm{x}_{n})=g\left(g(\bm{x}_{i-1};\bm{x}_{i});\bm{x}_{i+1}\right),\quad\text{where}\quad\bm{x}_{i},\bm{x}_{i+1}\in A.$ (1)
Here, the input sequence $X=(\bm{x}_{1},\ldots,\bm{x}_{n})$ comprises $n$ tokens, each of dimension $d$. An anchor set $A=\{\bm{a}_{1},\bm{a}_{2},\ldots,\bm{a}_{J}\}$ is designated, with each token $\bm{a}_{k}\in A$ corresponds to a function $g(\bm{x};\bm{a}_{k})$. In each $X$, one and only one pair of two consecutive elements belong to $A$, such as $\bm{x}_{i},\bm{x}_{i+1}\in A$. To simplify notation, we denote the two-anchor composite function as $f(\bm{x}_{i-1};\bm{x}_{i},\bm{x}_{i+1})$ to emphasize the composite anchor pair $(\bm{x}_{i},\bm{x}_{i+1})$ and key item $\bm{x}_{i-1}$.
For example, consider a simple case where $d=1$ and the anchor set $A=\{1,2,3,4\}$. Each anchor token corresponds to a specific function:
@@
Given $X=(23,1,2,43,46,74,44,72)$, the two-anchor composite function would output:
$f(X)=f(23;1,2)=g(g(23;1);2)=(23+5)+1=29.$ (3)
@@ 3.2 Data Generation
In this work, we construct the input dataset using four anchors (i.e., 1, 2, 3, 4) and 80 key items (i.e., 20-99). The four anchors comprise 16 composite anchor pairs. We select a subset or all of the composite anchor pairs to construct the training dataset based on the task requirements.
By default, i.e., with the inferential mapping defined below, the target is the output of the input data processed by the two-anchor composite function. Our definition of single-anchor function is consistent with the definition in Equ. (2).
3.3 Mapping Type of a Composite Anchor Pair
For a composite anchor pair ($a_{1}$, $a_{2}$), the type of its mapping $\mathcal{M}_{(a_{1},a_{2})}(\cdot)$ can be classified as follows.
Inferential mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is consistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=f(x;a_{1},a_{2})$.
Non-inferential mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is inconsistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)\neq f(x;a_{1},a_{2})$.
Symmetric mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is consistent with the mapping of its symmetric composite anchor pair ($a_{2}$, $a_{1}$), i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=\mathcal{M}_{(a_{2},a_{1})}(x)$.
We refer to a model as an inferential (non-inferential, symmetric) solution if it tends to output the mapping corresponding to the inferential (non-inferential, symmetric) mapping for the studied composite anchors.
@@ 3.4 Generalization
Generalization on data. Generalization on data primarily relies on the test set, which is divided according to key items. In this test data, all anchor pairs (i.e., the task) are seen in the training set.
Generalization on task. Generalization on task primarily depends on the unseen anchors, i.e., anchors that do not appear in the training set, with a designated target mapping.
@@ 3.5 Model Architecture and Basic Experimental Setups
where $X^{\mathrm{pos}}$ is the positional vector. For the $l$-th layer, the $Q,K,V$ are defined as:
@@
It is worth noting that the model architecture described above is for a single-head attention model.
@@ 4 Two Phases of Solutions for Composite Functions
In this section, we investigate the mechanisms of how a transformer network model fits the training data and its generalization.
@@
Experimental Setup. The anchor pair (4,3) is held out as an unseen pair during the training, while pair (3,4) is assigned as non-inferential mapping, i.e., we set the designated target mapping of ($\cdots$, $x$, 3, 4, $\cdots$) to $x-6$ (the inferential mapping is $x-10$), seen in the training set. The other 14 anchor pairs are set as inferential mappings seen in the training set. Three possible solutions for pair (4,3) may be learned after enough training: i) inferential solution based on the 14 anchor pairs with inferential mapping, ii) symmetric solution based on the anchor pair (3,4), iii) other non-inferential solutions. See illustration in Fig. 1(b).
@@
Figure 2: Phase diagram of generalization performance on unseen anchor (4, 3). (a) The model’s test accuracy based on the symmetric mapping. (b) The model’s test accuracy based on the inferential mapping. The abscissa represents the initialization rate $\gamma$, which corresponds to the standard deviation $(1/d_{\mathrm{in}})^{\gamma}$ of a normal distribution with a mean of 0 used for parameter initialization. The ordinate represents the depth of the transformer model. The shadow zones indicate the test accuracy on seen anchors is less than 90%.
An important and interesting question is: which solution is learned for unseen pair (4,3)? Through experiments with different parameter initialization scales and model depths, we find that the model use different mechanisms to fit training data, leading to different solutions for unseen pair (4,3). To see if the model learns the symmetric solution for the unseen anchor pair (4,3) during the training, we can plot the phase diagram of the model accuracy for the anchor pair (4,3) by assuming the ground truth as the symmetric mapping, against the parameter initialization scale and model depth. As shown in Fig. 2(a), the yellow zones with the large initialization scale (small $\gamma$) indicate $100\%$ test accuracy, i.e., the model learns the symmetric solution for anchor pair (4,3). Similarly, Fig. 2(b) assumes the ground truth as the inferential mapping. For small initialization scale (large $\gamma$), the model tends to learn the inferential solution for the anchor pair (4,3). When the initialization scale is too large and the model is deep (shadow zone), the model learns a non-inferential solution, which generalizes badly even on seen anchors.
We can conclude that the model tends to learn the inferential solution with small initialization while the symmetric solution with large initialization (but not very large). Note that this result is rather stable as we obtain these two phase diagrams over many trials as follows. For a specific random seed, models with different initializations and depths are trained using 9 learning rates (uniformly sampled in the interval [$1\times 10^{-4}$, $3\times 10^{-4}$]), and the highest accuracy for the inferential and symmetric mappings is selected. We use 3 random seeds for the experiments and take the average of the highest accuracy obtained from the three seeds.

### v2

# Initialization is Critical to Whether Transformers Fit Composite Functions by Inference or Memorizing

@@ Abstract
Transformers have shown impressive capabilities across various tasks, but their performance on compositional problems remains a topic of debate. In this work, we investigate the mechanisms of how transformers behave on unseen compositional tasks. We discover that the parameter initialization scale plays a critical role in determining whether the model learns inferential solutions, which capture the underlying compositional primitives, or symmetric solutions, which simply memorize mappings without understanding the compositional structure. By analyzing the information flow and vector representations within the model, we reveal the distinct mechanisms underlying these solution types. We further find that inferential solutions exhibit low complexity bias, which we hypothesize is a key factor enabling them to learn individual mappings for single anchors. Building upon the understanding of these mechanisms, we can predict the learning behavior of models with different initialization scales when faced with data of varying complexity. Our findings provide valuable insights into the role of initialization scale in shaping the type of solution learned by transformers and their ability to learn and generalize compositional tasks.
@@ 1 Introduction
Large-scale transformers (Vaswani et al., 2017) demonstrate unprecedented capabilities (Achiam et al., 2023; Brown et al., 2020; Choi et al., 2021; Teubner et al., 2023), even noted as "sparks of AGI" (Bubeck et al., 2023). However, their performance on compositional tasks, which are considered a key property of human cognition (Marcus, 2003; Smolensky et al., 2022), has long been the subject of intense debate, especially in terms of the out-of-distribution (OOD) generalization ability. This raises critical open questions about how to faithfully interpret transformers’ capabilities on compositional tasks. Can transformers learn the underlying compositional primitives within the data, or do they merely learn input-output mappings? When transformers produce incorrect outputs on compositional tasks, do their responses follow any discernible patterns or logic?
@@
In this work, we use anchor functions (Zhang et al., 2024), a framework for creating controlled synthetic data, to investigate how transformers generalize to unseen compositional tasks. In our setup, each sequence contains an anchor pair, a key item, and noise items unrelated to the output (right side of Fig. 1(a)). The single anchors are specific tokens (i.e., 1, 2, 3, 4), each corresponding to a particular arithmetic operation (left side of Fig. 1(a)). A composite function is defined as the sequential application of two single anchor functions, i.e., the corresponding arithmetic operation, to a key item. For example, given a key item $x$ and an anchor pair (1, 2), the composite function would output $(x+5)+1$, i.e., $x+6$. As shown in the middle part of Fig. 1(a), during training, 14 out of the 16 possible anchor pairs are assigned inferential mappings, meaning the target output is consistent with the result of applying the composite function. One pair (3, 4) is assigned a non-inferential mapping, where the target output deviates from the result of the composite function. The remaining pair (4, 3) is held out as an unseen task. After training, there are three possible solutions the model could learn for the unseen pair (4, 3): a symmetric solution matching the non-inferential seen pair (3, 4) (Mechanism 1 in Fig. 1(b)), an inferential solution (Mechanism 2 in Fig. 1(b)), or other non-inferential solutions. The central question we aim to answer is: which of these three types of solutions will the model learn for the unseen anchor pair (4, 3)?
@@
Figure 1: Experimental setup and possible solutions and mechanisms for the unseen anchor pair (4, 3). (a) Data generation: Left: The single anchors (i.e., 1, 2, 3, 4) correspond to specific arithmetic operations. Middle: During training, 14 out of the 16 possible anchor pairs are assigned inferential mappings, one pair (3, 4) is assigned a non-inferential mapping, and the remaining pair (4, 3) is held out as an unseen task (does not appear in the training). Right: The input sequences comprise an anchor pair, a key item preceding the anchor pair, and noise items unrelated to the target. The question mark indicates the output for the unseen anchor pair (4, 3), which depends on the learned solution. (b) Two potential mechanisms for the unseen anchor pair (4, 3): learning the symmetric structure (Mechanism 1) or composing the inferred single anchor mappings (Mechanism 2).
@@
Based on the above experimental setup, we find that the model exhibits two distinct phases of solutions depending on the parameter initialization scale with different model depths. With small initialization scales, the model tends to learn inferential solutions on the unseen anchor pair (4, 3). In contrast, with larger initialization scales, the model is more likely to learn symmetric solutions on anchor pair (4, 3) that simply match the output of its symmetric anchor pair (3, 4). These findings suggest that the initialization scale plays a crucial role in determining the type of solution learned by the model.
@@
Building upon our understanding of these mechanisms, we predict and experimentally validate that models with small initialization scales require more training steps to fit data with larger inferential complexity, defined as the number of anchor pair groups that do not satisfy the inferential mapping, while models with large initialization scales use similar training steps regardless of inferential complexity.
Our work highlights the crucial role of initialization scale in shaping the solutions learned by transformers on compositional tasks, enabling them to better capture compositional structures and generalize to unseen tasks. Furthermore, adapting the initialization scale to the specific task type could be promising: larger scales for memorization tasks such as text memorization, and smaller scales for inference tasks such as code generation. Unless otherwise specified, our main text primarily focuses on a single-head attention model to facilitate a clearer understanding of the underlying mechanisms. We further extend our experiments to multi-head attention models, such as GPT-2 (Radford et al., 2019), and verify that the insights drawn from the single-head model remain valid for more complex architectures.
@@ 2 Related Work
Challenges of transformers in compositional tasks. Recent advancements in large language models (LLMs) have showcased remarkable capabilities, often surpassing human performance (Fu et al., 2022; Wei et al., 2022). However, despite their impressive performance on single-step reasoning tasks (Srivastava et al., 2022), transformers struggle with multi-step compositional tasks and out-of-distribution (OOD) generalization (Csordás et al., 2021, 2022; Dziri et al., 2024; Hupkes et al., 2018; Lepori et al., 2023; Okawa et al., 2023; Yun et al., 2022). Ramesh et al. (Ramesh et al., 2023) show that transformers trained to directly compose capabilities struggle to generalize to OOD tasks with a synthetic task. Liu et al. (Liu et al., 2022) suggest that shallow transformers learn shortcuts during training, leading to poor OOD generalization. Numerous studies have explored various approaches to address these limitations, such as encouraging explicit reasoning step generation within a single generation (Wei et al., 2022), leveraging LLMs to generate reasoning steps iteratively (Creswell et al., 2022; Creswell and Shanahan, 2022). Despite these efforts, achieving complete mastery in compositional tasks remains a significant challenge for vanilla transformers. In order to clearly study the behaviors and internal mechanisms of language models, Zhang et al. Zhang et al. (2024) introduced anchor functions as benchmark functions for investigating transformer behavior. Our work builds on the anchor function setting to explore how different initialization scales affect model solutions and mechanisms.
Parameter initialization of neural networks. The parameter initialization of the network is important to determine the final fitting result of the network (Arora et al., 2019; Chizat and Bach, 2018; E et al., 2020; Jacot et al., 2018; Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2018; Sirignano and Spiliopoulos, 2020; Williams et al., 2019; Zhang et al., 2019). Luo et al. (Luo et al., 2021), Zhou et al. (Zhou et al., 2022) mainly identify the linear regime and the condensed regime for two-layer and three-layer wide ReLU NNs, respectively. A series of works suggests that the condensed networks are often accompanied by good generalization ability of the model (Zhang et al., 2021, 2022, 2023; Zhang and Xu, 2024). Recent studies have also investigated the impact of initialization on the training of LLMs (Huang et al., 2020; Liu et al., 2020; Trockman and Kolter, 2023; Wang et al., 2024; Zhang et al., 2019; Zhu et al., 2021). These works primarily focus on how the initialization scale affects the stability of the training process and plays a crucial role in ensuring smooth and effective training of LLMs. In our work, we find that different initialization scales can significantly influence a model’s ability to memorize and infer data on the compositional task, highlighting the profound impact of initialization on the final performance and underlying mechanisms of the trained models.
@@ 3 Definitions
We introduce a set of key definitions that will be used throughout the paper. For detailed explanations and illustrative examples to better understand these definitions, please refer to Appendix B.
@@ 3.1 Two-anchor composite function
A two-anchor composite function $f(X):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is defined as
$f(x_{1},\ldots,x_{n})=g\left(g(x_{i-1};x_{i});x_{i+1}\right),\quad\text{where}\quad x_{i},x_{i+1}\in A.$ (1)
Here, the input sequence $X=(x_{1},\ldots,x_{n})$ comprises $n$ tokens. An anchor set $A=\{a_{1},a_{2},\ldots,a_{J}\}$ is designated, with each token $a_{k}\in A$ corresponds to a function $g(x;a_{k})$. In each $X$, one and only one pair of two consecutive elements belong to $A$, such as $x_{i},x_{i+1}\in A$. We refer to the token immediately preceding the anchor pair as the key item. To simplify notation, we denote the two-anchor composite function as $f(x_{i-1};x_{i},x_{i+1})$ to emphasize the anchor pair $(x_{i},x_{i+1})$ and key item $x_{i-1}$.
In this work, we set the anchor set $A=\{1,2,3,4\}$. Each anchor token corresponds to a specific function:
@@ 3.2 Data Generation
In this work, we construct the input dataset using four anchors (i.e., 1, 2, 3, 4) and items sampled from 20-99. Each sequence includes an anchor pair, a key item (the item immediately preceding the anchor pair), and noise items. The noise items are unrelated to the target. The four anchors form 16 anchor pairs, and we select a subset or all of these pairs to construct the training dataset based on the task requirements.
By default, the target is the output of the input data processed by the two-anchor composite function, i.e., the inferential mapping defined below. We divide the training data and test data based on the value of the key item. The specific division method can be found in Appendix C.
3.3 Mapping Type of an anchor pair
For an anchor pair ($a_{1}$, $a_{2}$), the type of its mapping $\mathcal{M}_{(a_{1},a_{2})}(\cdot)$ can be classified as follows.
Inferential mapping. The designated target mapping of an anchor pair ($a_{1}$, $a_{2}$) is consistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=f(x;a_{1},a_{2})$.
Non-inferential mapping. The designated target mapping of an anchor pair ($a_{1}$, $a_{2}$) is inconsistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)\neq f(x;a_{1},a_{2})$.
Symmetric mapping. The designated target mapping of an anchor pair ($a_{1}$, $a_{2}$) is consistent with the mapping of its symmetric anchor pair ($a_{2}$, $a_{1}$), i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=\mathcal{M}_{(a_{2},a_{1})}(x)$.
We refer to a model as an inferential (non-inferential, symmetric) solution on an anchor pair if it tends to output the mapping corresponding to the inferential (non-inferential, symmetric) mapping for the studied pair.
@@ 3.4 Generalization
Generalization on data. Generalization on data relies on the test set (defined in Appendix C). In this test data, all anchor pairs (i.e., the task) are seen in the training set.
Generalization on task. Generalization on task depends on the unseen anchors, i.e., anchors that do not appear in the training set, with a designated target mapping.
@@ 3.5 Model Architecture and Basic Experimental Setups
To enable a more focused analysis of the underlying mechanisms, the following sections will only introduce the architecture of the single-head attention model.
@@
where $X^{\mathrm{pos}}$ is a trainable positional vector. For the $l$-th layer, the $Q,K,V$ are defined as:
@@ 4 Two Phases of Solutions for Composite Functions
In this section, we investigate the mechanisms of how a transformer learns the compositional tasks, especially on the OOD tasks.
@@
Experimental Setup. For the 16 anchor pairs, the anchor pair (4, 3) is held out as an unseen pair during the training, while pair (3, 4) is assigned as non-inferential mapping, i.e., we set the designated target mapping $\mathcal{M}_{(3,4)}(x)=x-6$ (the inferential mapping is $x-10$), seen in the training set. The other 14 anchor pairs are set as inferential mappings seen in the training set. Three possible solutions for pair (4, 3) may be learned after enough training: i) inferential solution based on the 14 anchor pairs with inferential mapping, ii) symmetric solution based on the anchor pair (3, 4), iii) other non-inferential solutions. See illustration in Fig. 1(b).
@@
Figure 2: (a,b) Phase diagram of generalization performance on the unseen anchor (4, 3). (a) The model’s test accuracy based on the symmetric mapping. (b) The model’s test accuracy based on the inferential mapping. The abscissa represents the initialization rate $\gamma$, which corresponds to the standard deviation $(1/d_{\mathrm{in}})^{\gamma}$ of a normal distribution with a mean of 0 used for parameter initialization. The ordinate represents the depth of the transformer model. The shadow zones indicate the test accuracy on seen anchors is less than 90%. (c) Comparison of accuracy on the unseen anchor (4, 3) for both the inferential and symmetric solutions across different initialization rates $\gamma$ on GPT-2-SMALL. The error bars represent the standard deviation across 4-time runs.
An important and interesting question is: which solution is learned for unseen pair (4, 3)? Experiments show that different initialization scales lead to different solutions for the unseen pair (4, 3) with different model depths. Fig. 2(a) shows the phase diagram assuming the symmetric mapping ($\mathcal{M}_{(4,3)}(x)=x-6$) as ground truth, with initialization scale on the abscissa and model depth on the ordinate. A large initialization scale (small $\gamma$) leads to the symmetric solution, indicated by yellow zones with nearly 100% accuracy. Fig. 2(b) assumes the inferential mapping ($\mathcal{M}_{(4,3)}(x)=x-10$) as ground truth, showing that a small initialization scale (large $\gamma$) favors the inferential solution. Larger initialization scales with deep models result in poor generalization even on seen anchors (shadow zone). This pattern was consistent across trials: for each random seed, models with different initializations and depths were trained using 9 learning rates (uniformly sampled in the interval [$1\times 10^{-4}$, $3\times 10^{-4}$]), and the highest accuracy for each of the two mappings was selected from these rates. The highest accuracies were then averaged over 3 random seeds. We conclude that small initializations favor the inferential solution, while large (but not very large) initializations favor the symmetric solution.
To validate the generality of these findings, we extend our experiments to the multi-head GPT-2-SMALL model. As shown in Fig. 2(c), the model exhibits similar generalization performance trends on the unseen pair (4, 3) across different initialization rates $\gamma$. The range of $\gamma$ variation is smaller in this setup, likely due to the difference in model parameter scale between the GPT-2 and the transformer models used in the previous experiments.
We also find that increasing the learning rate and weight decay coefficient within a certain range is helpful for the transformer to learn the inferential solution on the unseen anchor. Please refer to Appendix E for experimental phenomena and detailed discussion.